# DEPTH: Discourse Education through Pre-Training Hierarchically

### v1

@@ Appendix E GLUE Tables
In this section we show the full results from our downstream experiments on GLUE tasks. Table 3 is reflected in the top row (FS) of Figure 4, while Table 4 is reflected in the bottom row (CPT) of figure 4. Consistent with our hypothesis, we found that both DEPTH and T5 improve across downstream tasks as a function of the pre-training steps they’ve taken. While DEPTH outperforms T5 across all tasks is the FS setting, it did not reach the scores of Raffel et al. [45]’s T5 model. In the CPT setting, T5 and DEPTH perform quite comparably. In fact, in the penultimate checkpoint (128k) we found that DEPTH outperformed T5 on all tasks except for CoLA.
@@
Appendix F DiscoEval Tables
@@
Table 5: DiscoEval Downstream Full Training (FS) Results
@@ Appendix G NI Results
Taken as a whole, while DEPTH performs well on NLU tasks (e.g, DC and SP from DiscoEval), when applied with limited pre-training and no supervised fine-tuning [60], its performance may be weaker than other encoder-decoders with respect to downstream tasks involving long-form text generation.

### v2

@@ Appendix F GLUE results
In this section we show the full results from our downstream experiments on GLUE tasks. Table 5 is reflected in the top row (FS) of Figure 4, while Table 6 is reflected in the bottom row (CPT) of figure 4. Consistent with our hypothesis, we found that both DEPTH and T5 improve across downstream tasks as a function of the pre-training steps they’ve taken. While DEPTH outperforms T5 across all tasks is the FS setting, it did not reach the scores of Raffel et al. (2020)’s T5 model. In the CPT setting, T5 and DEPTH perform quite comparably. In fact, in the penultimate checkpoint (128k) we found that DEPTH outperformed T5 on all tasks except for CoLA.
@@
Appendix G DiscoEval results
@@
Table 7: DiscoEval Downstream Full Training (FS) Results